# Not flexible enough? Impacts of electric carsharing on a power sector with variable renewables11footnotemark: 1

### v1

@@ Abstract
Electrifying the car fleet is a major strategy for mitigating greenhouse gas emissions in the transport sector. However, electrification alone will not solve all the negative externalities associated with cars. In light of other problems such as congestion, noise and street space as well as concerns about the use of mineral resources for battery electric cars, reducing the size of the car fleet would be beneficial, particularly in cities. Carsharing could offer a way to reconcile current car usage habits with a reduction in the size of the car fleet. However, it could also reduce the potential of electric cars to align their grid interactions with variable renewable electricity generation. We investigate how electric carsharing may impact the power sector in the future. To do so, we combine three open-source quantitative methods, including sequence clustering of car travel diaries, a probabilistic tool to generate synthetic electric vehicle time series, and an optimization model of the power sector. For 2030 scenarios of Germany with a renewable share of at least 80%, we show that switching to electric carsharing only moderately increases power sector costs. In our main setting with a mixed fleet including shared and private electric vehicles, carsharing increases yearly power sector costs by less than 100 euros per substituted private electric car. This cost effect increases with the share of shared cars in the overall fleet and is largest under the assumption of bidirectional charging. It is mitigated when other sources of flexibility for the power sector are considered, such as a flexible green hydrogen industrial demand. Carsharing further causes a shift from wind power to solar PV in the optimal capacity mix, and may also trigger additional investments in stationary electricity storage, especially in flexibility-constrained settings. Overall, we find that shared electric cars still have the potential to be operated largely in line with variable renewable electricity generation. We conclude that electric carsharing is unlikely to cause much damage to the power sector, but could bring various other benefits, which may outweigh power sector cost increases.
@@ 1 Introduction
Electrifying the passenger vehicle fleet is a major strategy for reducing greenhouse gas emissions in the passenger transport sector (noauthor_global_2023 (19, 4)). This is because battery electric vehicles (BEVs) often have lower life-cycle emissions than internal combustion engine vehicles (ICEVs) already today, and potentially much more so in the future (sacchi_when_2022 (34, 5, 22, 18)). As policymakers increasingly support the electric vehicle industry, the electrification of passenger cars is also becoming more and more widespread. This trend is expected to continue and accelerate as bans on the sale of new non-zero emission cars are announced in various parts of the world. In the European Union, the ban will take effect from 2035. This will also be the case in some states of the United States such as California. In China, at least half of all new cars sold after 2035 will have to be electric, fuel cell or plug-in hybrid.
@@
At the same time, there is a growing interest in carsharing, especially in urban areas (munzel2020explaining (32, 12, 23)). By reducing the overall fleet size, carsharing has the potential to increase available street space in cities (creutzig2020fair (6, 14)). Carsharing and car electrification are interrelated, as the former could also foster consumer uptake of electric cars. Shared cars could thus play a role in bringing the electrification of the car fleet to scale (mccollum2018interaction (28)). However, while electric carsharing offers numerous benefits, it also poses some challenges for the power sector. One challenge is the potential loss of flexibility. Privately-owned BEVs with system-oriented charging capabilities can provide flexibility to the power sector. This facilitates the integration of variable renewable energy sources such as solar and wind power (liu2024 (24, 44, 41, 17)), and can help to shave peak loads (xu2018planning (45)). This way, BEVs can also reduce stationary electricity storage needs (xu2023storage (43)) and respective use of mineral resources (deetman2021 (8)). Electric carsharing may reduce the flexibility potential of BEVs due to their increased driving frequency, which translates into a lower grid availability, as well as lower aggregate battery capacity of the fleet.
@@
Taking into account both the trends of electrification of passenger transport on the one hand and the development of sharing options on the other, we investigate the impacts of a switch from private BEVs to electric carsharing on a power sector dominated by variable renewable energy sources. In our study, carsharing refers to a free-floating fleet of shared vehicles that can be rented on a per-trip basis. Carsharing differs from ride-sharing or share-pooling in that individuals do not share a vehicle during a journey. In Germany, carsharing has been in use for longer than ride-pooling and is therefore more established (burghard2022sharing (3)). Carsharing might contribute to reducing car ownership (jochem_does_2020 (20, 31)). Therefore, carsharing can also be relevant in the transport decarbonization strategy, as a reduction in the size of the car fleet may be desirable beyond the mere electrification of the passenger road transport sector in order to achieve decarbonization targets (dirnaichner_life-cycle_2022 (10, 29)). Case studies have shown that carsharing can also help to reduce other negative externalities, and, for instance, increase the availability of public space in central city areas (diana_spatial_2022 (9)). Finally, carsharing could help mitigate negative impacts of cars on marine and freshwater ecotoxicity as well as on mineral resource scarcity (vilaca_life_2022 (42)). Hence, electric carsharing could also help tackle the potential trade-off between decarbonizing transport and the need for critical metals (zhang_trade-off_2023 (47, 46)).
@@
In addition to the challenge of electrifying the passenger car fleet, meeting decarbonization targets also requires an increasing use of electricity generated from renewable energy sources. In most countries, the renewable energy sources with the largest potential are wind and solar power, which have variable generation patterns (lopezprol2021 (26)). This requires future power sectors to be more flexible. Various options for temporal and spatial flexibility exist, including various types of energy storage (schill2020 (35)). While electrifying passenger road transport increases electricity demand, and thus the need for integrating variable renewables, it may also increase the power sector flexibility if BEVs are charged smartly or bidirectionally using vehicle-to-grid (V2G) technologies (syla2024 (40, 27, 36, 25)). In particular, it has been found that BEVs may help balance daily variations of solar power (brown2018synergies (2)). One important factor enabling this flexibility is the fact that many privately-owned vehicles remain idle for extended periods of time. Assuming sufficient grid connection, electric vehicles could charge during times of abundant renewable power and, if bidirectional charging is possible, feed electricity back into the grid during times of scarcity. The actual flexibility potential that BEV can provide to the power sector depends on various factors, such as driving profiles and plug-in behaviours (gschwendtner_impact_2023 (16)). Furthermore, the overall battery capacity of the BEV fleet that is equipped with smart or bidirectional charging is also a significant factor.
@@
A few analyses focus specifically on the potential of carsharing fleets for grid congestion management (brinkel2022grid (1)) or take the perspective of operators of shared BEV fleets (kahlen2018electric (21)) or the end users’ perspective to investigate whether carsharing users are willing to offer temporal flexibility when using shared BEVs (suel2024vehicle (39)). We aim to address this research gap with a quantitative, model-based study of the power sector effects related to switching from privately-owned BEVs to electric carsharing. We do so for 2030 scenarios of the German power sector with a renewable energy share of at least 80%. We use a combination of three open-source quantitative methods: (i) sequence clustering techniques on car travel diaries from a national German mobility survey, (ii) emobpy, a probabilistic tool for generating synthetic BEV time series, and (iii) DIETER, a capacity expansion model of the power sector. For clarity, we assume no changes in demand for car mobility, and consider three different BEV charging strategies which apply to the whole vehicle fleet.
@@ 2 A combination of three quantitative methods
Our methodological approach is based on a three-step procedure (Figure 1). First, probability distributions for car mobility are computed based on empirical data. These distributions are generated for both private and shared BEVs, assuming a constant substitution rate. Secondly, the open-source tool emobpy is used to generate time series for the electric consumption and charging availability of BEVs in an hourly resolution. Thirdly, these time series serve as input for a linear optimization open-source power sector model called DIETER. The first two steps are exclusively carried out using Python, while the third step also involves GAMS programming. All codes are available on the dedicated GitLab repository. [^2] The repository can be accessed at https://gitlab.com/diw-evu/electric-carsharing.
@@ 2.1 Sequence clustering
To generate mobility probability distributions, we use travel diaries from the representative survey Mobilität in Deutschland (MiD) (eggs2018mobilitat (11)). The travel diaries are reshuffled into a sequence format, and a clustering algorithm is applied to identify distinct groups of car usage.
@@ 2.1.1 Travel diaries
GPS tracking and big data methods can provide access to highly detailed time series with high geographical and temporal resolution for existing car users. However, electric cars and carsharing are still in the early stages of development. Therefore, generalizing the behaviour of early adopters to a country-wide setting could lead to flawed conclusions. This study departs from existing empirical data on electric car or carsharing use and instead uses travel diaries from a representative survey, which allows to consider geographical components that are relevant in this context, such as the level of urbanization. In addition, travel diaries provide detailed information on car use at a high temporal resolution.
This study uses the B.1 version of the 2017 edition of the MiD, commissioned by the German Federal Ministry for Transport and Digital Infrastructure. [^3] Bundesministerium für Digitales und Verkehr (BMDV). The data collection took place in Germany from May 2016 to September 2017. A total of 316,361 individuals from 156,420 households were surveyed, representing the German population at the time (around 41 million households and 82 million individuals). For the analysis, we only considered car trips with available departure and arrival times. We excluded passenger trips to avoid double counting, as most car trips taken as a passenger are with another household member as the driver. We also excluded car trips undertaken during one’s professional activity. Individuals whose diary implied overlapping trips, i.e. at least two trips taking place at least partially at the same time, and trips that span two days were also removed. The final database comprises 355,136 car trips.
@@ 2.1.3 Clustering
We implement the Hierarchical Agglomerative Clustering (HAC) algorithm using the Levenshtein distance as a distance measure and Ward’s method as the linkage criterion. The clustering is performed on subdatasets that partition the overall sequence dataset by location type: metropolis, big city, central city/medium-sized city, urban area, and small town/village area. [^4] The location typology corresponds to the variable RegioStatR Gem5 provided in the MiD dataset. Additional information is available on the BMDV website at this link. The final number of clusters is selected using dendrograms (Figure 7).
The computed descriptive statistics for the clusters (Table 3) indicate that car usage can be characterised by the daily distance travelled by car, which correlates well with the average trip distance and duration. For example, individuals in cluster 6 travel approximately 11 kilometers by car, with an average trip distance of 5.5 kilometers and an average trip duration of 14 minutes. In contrast, individuals in cluster 4 travel an average of 50 kilometers, with an average trip distance of 18.5 kilometers and an average trip duration of 34.2 minutes. To visualize differences in car use, we refer to sequence index plots (Figure 8).
This applies to all locations, even though there may be some differences in the characteristics of the clusters across locations. For example, in the cluster of users who drive short distances overall during the day, the total distance travelled as well as the average trip distance are greater in less urbanized areas. Conversely, the average trip duration is bigger in more urbanized areas, which is in line with congestion being more prevalent in big cities than rural areas.
@@ 2.2 Synthetic BEV time series
Using the open-source tool emobpy, [^5] For more detailed information on emobpy, we refer to the dedicated publication (gaete-morales_open_2021 (13)) and the appendix (B). we generate synthetic BEV time series for private and shared BEVs based on clustering results. Synthetic time series are computed at the cluster level for each location at a 5-minute resolution. In all computations, we use the default values for the Volkswagen ID.3 model. For shared BEVs across all clusters and private BEVs from the cluster with longest trips, we assume a battery capacity of 100 kWh, compared to 58 kWh in other cases. Probability distributions for charging station availability can be found in Table 5. The charging strategy used to compute the grid electricity time series is balanced.
@@ 2.2.1 Vehicle mobility
Assuming a substitution rate of five between private cars and shared cars, we derive probability distributions for shared BEVs based on the probability distributions for private BEVs. The distribution of number of trips for shared BEVs is the same as for private BEVs, but centered around a new mean (see B). We assume that the distribution of trip departure times remains the same as in the private case, except that we condition only on the trip destination. For the joint distribution of trip duration and distance, we also assume the same as in the private case and remove the conditioning on the total number of trips (Table 4). Thus, this modelling approach is based on the assumption that carsharing does not induce a change in the number or type of car trips made.
@@
2.3 Power system modelling
We apply the open-source power sector model DIETER (Dispatch and Investment Evaluation Tool with Endogenous Renewables). As a capacity expansion model, it minimizes fixed and variable power sector costs under various feasibility and policy constraints. The model is solved for a full year in an hourly resolution (8760 consecutive hours). Its main focus is the integration of variable renewable energy sources (vRES) by means of various options for power sector flexibility. We apply the model to a 2030 scenario of Germany with an 80% renewable share. For clarity and generalizability, we abstract from the European interconnection and model Germany as an electric island, assuming perfect expansion of the transmission and distribution grids within the country. Sector coupling with the heating sector is modelled assuming a stock of 7.5 million heat pumps, which are assumed to be operated inflexibly, i.e. to provide heat at the time of demand. [^6] For more details, see roth2023flexible (33). Using around 52 TWh of electricity, these heat pumps generate around 143 TWh of heat, corresponding to around twenty percent of the annual space heating demand from residential buildings in Germany. In additional model runs, we also include a 30 TWh flat annual demand for green hydrogen, which corresponds to an additional electricity demand of around 42 TWh.

### v2

# Impacts of electric carsharing on a power sector with variable renewables11footnotemark: 1

@@ Abstract
Electrifying the car fleet is a major strategy for mitigating emissions in the transport sector. As electrification cannot solve all negative externalities associated with cars, reducing the size of the car fleet would be beneficial. Electric carsharing could allow to reconcile current car usage habits with a smaller fleet, but this may reduce the potential of electric cars to align their grid interactions with variable renewable electricity generation. We investigate how electric carsharing may impact the power sector, combining three methods: sequence clustering of car travel diaries, generation of synthetic electric vehicle time series, and power sector modelling. We show that switching to electric carsharing only moderately increases power sector costs, less than 110 euros per substituted car in our main setting. This effect is largest with bidirectional charging. We conclude that the power sector interactions of shared electric car fleets could still be aligned with variable renewable electricity generation.
@@ 1 Introduction
Electrifying the vehicle fleet is a major strategy for reducing greenhouse gas (GHG) emissions in the passenger road transport sector, which represents the largest source of GHG emissions in the transport sector ([noauthor_global_2023, ipcc2023]). This is because battery electric vehicles (BEVs) often have lower life-cycle emissions than internal combustion engine vehicles (ICEVs) already today, and potentially much more so in the future ([sacchi_when_2022, cox_life_2020, knobloch_net_2020, hoekstra2019]). As policymakers increasingly support the electric vehicle industry, the electrification of passenger cars is also becoming more and more widespread. This trend is expected to continue and accelerate as bans on the sale of new non-zero emission cars are announced in various parts of the world. In the European Union, the ban will take effect in 2035. This will also be the case in some states of the United States such as California. In China, at least half of all new cars sold after 2035 will have to be electric, fuel cell or plug-in hybrid.
@@
At the same time, there is a growing interest in carsharing, especially in urban areas ([creutzig2024, fulton2018, kortum2016free]), where cars are responsible for environmental externalities ([leroutier2023tackling]). By reducing the overall fleet size, carsharing has the potential to increase available street space in cities ([creutzig2020fair, glotz2016reclaim]). Carsharing and car electrification are interrelated, as the former could also foster consumer uptake of electric cars. Shared cars could thus play a role in bringing the electrification of the car fleet to scale ([hoerler2021carsharing, shaheen2020zero, schluter2019car, mccollum2018interaction]). Besides, as light-duty vehicles represent the main mode of transport for private citizens ([ipcc2023]), carsharing can be perceived as a more appealing transport mode than other mobility options such as public transport or active mobility, as it could require smaller behavioral adaptation compared to a so-called “modal shift”. However, while electric carsharing offers numerous benefits, it also poses some challenges to the power sector. One challenge is the potential loss of flexibility. privately owned BEVs with system-oriented charging capabilities can provide flexibility to the power sector. This facilitates the integration of variable renewable energy sources (vRES) such as solar and wind power ([liu2024, xu2023, taljegard2019impact, hanemann2017effects]) and can help to shave peak loads ([xu2018planning]). This way, BEVs can also reduce stationary electricity storage needs ([xu2023storage]) and respective use of mineral resources ([deetman2021]). Electric carsharing may reduce the flexibility potential of BEVs due to their increased driving frequency, which translates into a lower grid availability, as well as to the lower aggregate battery capacity of the fleet.
@@
Taking into account both the trend of electrification of passenger transport on the one hand and the development of sharing options on the other, we investigate the impacts of a switch from private BEVs to electric carsharing on a power sector dominated by variable renewable energy sources. In our study, carsharing refers to a free-floating fleet of shared vehicles that can be rented on a per-trip basis. Carsharing differs from ride-sharing or share-pooling in that individuals do not share a vehicle during a journey. In Germany, carsharing has been in use for longer than ride-pooling and is therefore more established ([burghard2022sharing]). Carsharing might contribute to reducing car ownership ([jochem_does_2020, mounce_potential_2019]) and hence the car fleet size ([spieser2014toward]). Therefore, carsharing can also be relevant in the transport decarbonization strategy, as a reduction in the size of the car fleet may be desirable beyond the mere electrification of the passenger road transport sector in order to achieve decarbonization targets ([dirnaichner_life-cycle_2022, milovanoff2020electrification]). Case studies have shown that carsharing can also help reduce other negative externalities and, for instance, increase the availability of public space in central city areas ([diana_spatial_2022]). Finally, carsharing could help mitigate negative impacts of cars on marine and freshwater ecotoxicity as well as on mineral resource scarcity ([vilaca_life_2022]). Hence, electric carsharing could also help tackle the potential trade-off between decarbonizing transport and the need for critical metals ([zhang_trade-off_2023, zeng2022]).
@@
In addition to the challenge of electrifying the passenger car fleet, meeting decarbonization targets also requires an increasing use of electricity generated from renewable energy sources (RES). In most countries, the renewable energy sources with the largest potential are wind and solar power, which have variable generation patterns ([lopezprol2021]). This requires future power sectors to be more flexible. Various options for temporal and spatial flexibility exist, including various types of energy storage ([schill2020]). While electrifying passenger road transport increases electricity demand, and thus the need for integrating variable renewables, it may also increase the power sector flexibility if BEVs are charged smartly or bidirectionally using vehicle-to-grid (V2G) technologies ([syla2024, luca2018impact, schill2015power, loisel2014large]). In particular, it has been found that BEVs may help balance daily variations of solar power ([brown2018synergies]). One important factor enabling this flexibility is the fact that many privately owned vehicles remain idle for extended periods of time. Assuming sufficient grid connection, electric vehicles could charge during times of abundant renewable power and, if bidirectional charging is possible, feed electricity back into the grid during times of scarcity. The actual flexibility potential that BEVs can provide to the power sector depends on various factors, such as driving profiles and plug-in behaviours ([gschwendtner_impact_2023]). Furthermore, the overall battery capacity of the BEV fleet that is equipped with smart or bidirectional charging is also a significant factor.
@@
A few analyses focus specifically on the potential of carsharing fleets for grid congestion management ([brinkel2022grid]) or take the perspective of operators of shared BEV fleets ([wiedemann2024vehicle, kahlen2018electric]) or the end users’ perspective to investigate whether carsharing users are willing to offer temporal flexibility when using shared BEVs ([suel2024vehicle]). To the best of our knowledge, the effects on the power sector’s optimal investment and dispatch related to switching from privately owned BEVs to electric carsharing have not been studied yet. We aim to address this research gap with a quantitative model-based study. We do so for 2030 scenarios of the German power sector with renewable energy shares of around 85%. We use a combination of three open-source quantitative methods: (i) sequence clustering techniques on car travel diaries from a national German mobility survey, (ii) emobpy, a probabilistic tool for generating synthetic BEV time series, and (iii) DIETER, a capacity expansion and dispatch model of the power sector. For clarity, we assume no changes in demand for car mobility, and consider three different BEV charging strategies which apply to the whole vehicle fleet.
We show that a switch to electric carsharing decreases the flexibility potential of BEVs and thus increases power sector costs. This cost increase however remains moderate in most settings analyzed, and remains below 110 euros per substituted car and year in our main scenario. Costs effects of carsharing are driven by changes in the optimal power plant portfolio and its hourly use. The nature and magnitude of impacts of switching to carsharing depend on the charging strategy considered. Costs increase the most when all electric vehicles are assumed to be charged bidirectionally, i.e. make full use of their flexibility potential. A switch to carsharing still appears to be compatible with a large share of renewable energy in the electricity mix of about 85%. Overall, our research demonstrates that future BEV fleets that include relevant amounts of shared cars can still be operated largely in line with variable renewable energy sources and contribute to their system integration.
@@ 2 A combination of three quantitative methods
Our methodological approach is based on a three-step procedure (Figure 1). First, probability distributions for car mobility are computed based on empirical data. These distributions are generated for both private and shared BEVs, assuming a constant substitution rate. Secondly, the open-source tool emobpy is used to generate time series for the electric consumption and grid availability of BEVs at an hourly resolution. Thirdly, these time series serve as input for a linear optimisation open-source power sector model called DIETER. The first two steps are carried out using Python exclusively, while the third step also involves GAMS programming.
@@ 2.1 Sequence clustering
To generate mobility probability distributions, we use travel diaries from the representative survey Mobilität in Deutschland (short MiD, see [eggs2018mobilitat]). The travel diaries are reshuffled into a sequence format, and a clustering algorithm is applied to identify distinct groups of car usage.
@@ 2.1.1 Travel diaries
GPS tracking and big data methods can provide access to highly detailed time series with high geographical and temporal resolution for existing car users. However, electric cars and carsharing are still in the early stages of development. Therefore, generalizing the behaviour of early adopters to a country-wide setting could lead to flawed conclusions ([yeh2022improving]). This study departs from existing empirical data on electric car or carsharing use and instead uses travel diaries from a representative survey, which allows to consider geographical components that are relevant in this context, such as the level of urbanization. In addition, travel diaries provide detailed information on car use at a high temporal resolution.
This study uses the B1 version of the 2017 edition of the MiD, commissioned by the German Federal Ministry for Transport and Digital Infrastructure (Bundesministerium für Digitales und Verkehr, BMDV). The data collection took place in Germany from May 2016 to September 2017. A total of 316,361 individuals from 156,420 households were surveyed, representing the German population at the time (around 41 million households and 82 million individuals). For the analysis, we only consider car trips with available departure and arrival times. We exclude passenger trips to avoid double counting, as most car trips taken as a passenger are with another household member as the driver. We also exclude car trips undertaken during one’s professional activity. Individuals whose diary implied overlapping trips, i.e. at least two trips taking place at least partially at the same time, and trips that span two days were also removed. The final database comprises 355,136 car trips.
@@ 2.1.3 Clustering
We implement the Hierarchical Agglomerative Clustering (HAC) algorithm using the Levenshtein distance as a distance measure and Ward’s method as the linkage criterion. The clustering is performed on subdatasets that partition the overall sequence dataset by location type (metropolis, big city, middle-size city, small city, and rural area). The location typology corresponds to the variable RegioStatR Gem5 provided in the MiD dataset. Additional information is available on the BMDV website at this link. The final number of clusters is selected using dendrograms (Figure 7).
The computed descriptive statistics for the clusters (Table 3) indicate that car usage can be characterised by the daily distance travelled by car, which correlates well with the average trip distance and duration. For example, individuals in cluster 6 travel approximately 11 kilometers by car, with an average trip distance of 5.5 kilometers and an average trip duration of 14 minutes. In contrast, individuals in cluster 4 travel an average of 50 kilometers, with an average trip distance of 18.5 kilometers and an average trip duration of 34.2 minutes. To visualise differences in car use, we refer to sequence index plots (Figure 8).
This applies to all locations, even though there may be some differences in the characteristics of the clusters across locations. For example, in the cluster of users who drive short distances overall during the day, the total distance travelled as well as the average trip distance are greater in less urbanized areas. Conversely, the average trip duration is bigger in more urbanised areas, which is in line with congestion being more prevalent in big cities than rural areas.
@@ 2.2 Synthetic BEV time series
Using the open-source tool emobpy, we generate synthetic BEV time series for private and shared BEVs based on clustering results. Synthetic time series are computed at the cluster level for each location at a 5 -minute resolution. In all computations, we use the default values for the Volkswagen ID.3 model. For shared BEVs across all clusters and private BEVs from the cluster with longest trips, we assume a battery capacity of 100 kWh, compared to 58 kWh for other private vehicles. As shared BEVs must be able to make more trips per day, we assume that such vehicles would be equipped with bigger batteries than privately owned cars that are used only by one household. This battery capacity is slightly above the battery capacity of models currently offered by carsharing operators. Yet, it appears plausible that future models will be equipped with bigger batteries, as the battery manufacturing costs keep decreasing. Already today, carsharing operators provide models with above-average batteries, such as the Tesla model Y, the Tesla model 3 or the VW ID.4, with battery capacities of about 80 kWh. Probability distributions for charging station availability can be found in Table 5. The charging strategy used to compute the grid electricity time series is balanced. For more detailed information on emobpy, we refer to the dedicated publication ([gaete-morales_open_2021]) and the appendix (B).
@@ 2.2.1 Vehicle mobility
Assuming a substitution rate of about five between private cars and shared cars, we derive probability distributions for shared BEVs based on the probability distributions for private BEVs. This substitution rate can be seen as a rather conservative assumption since previous research has shown that free-floating carsharing could replace between 8 and 20 private cars ([jochem_does_2020]). Yet, the conditions under which carsharing would substitute private car ownership are still debated ([silvestri2021contribution]). In this regard, assuming a conservative substitution rate of about five by 2030 seems reasonable. The distribution of number of trips for shared BEVs is the same as for private BEVs, but centered around a new mean (see B). We assume that the distribution of trip departure times remains the same as in the private case, except that we condition only on the trip destination. For the joint distribution of trip duration and distance, we also assume the same as in the private case and remove the conditioning on the total number of trips (Table 4). Thus, this modelling approach is based on the assumption that carsharing does not induce a change in the number or type of car trips made.
@@
2.3 Power sector modelling
We apply the open-source power sector model DIETER (Dispatch and Investment Evaluation Tool with Endogenous Renewables). As a capacity expansion model, it minimizes fixed and variable power sector costs under various feasibility and policy constraints. The model is solved for a full year at an hourly resolution (8760 consecutive hours). Its main focus is the integration of variable renewable energy sources by means of various options for power sector flexibility. We apply the model to a 2030 scenario of Germany with a renewable share of around 85%. For clarity and generalizability, we abstract from the European interconnection and model Germany as an electric island, assuming perfect expansion of the transmission and distribution grids within the country. Sector coupling with the heating sector is modelled assuming a stock of 7.5 million heat pumps, which are assumed to be operated inflexibly, i.e. to provide heat at the time of demand. The electricity consumption time series of heat pumps has been derived from [roth2023flexible]. Using around 52 TWh of electricity, these heat pumps generate around 143 TWh of heat, corresponding to abou twenty percent of the annual space heating demand from residential buildings in Germany. In additional model runs, we also include a 30 TWh flat annual demand for green hydrogen, which corresponds to an additional electricity demand of around 42 TWh.